# MSFMamba: Multi-Scale Feature Fusion State Space Model for Multi-Source Remote Sensing Image Classification

### v1

@@ Abstract
In multi-source remote sensing image classification field, remarkable progress has been made by convolutional neural network and Transformer. However, existing methods are still limited due to the inherent local reductive bias. Recently, Mamba-based methods built upon the State Space Model have shown great potential for long-range dependency modeling with linear complexity, but it has rarely been explored for the multi-source remote sensing image classification task. To this end, we propose Multi-Scale Feature Fusion Mamba (MSFMamba) network for hyperspectral image (HSI) and LiDAR/SAR data joint classification. Specifically, MSFMamba mainly comprises three parts: Multi-Scale Spatial Mamba (MSpa-Mamba) block, Spectral Mamba (Spe-Mamba) block, and Fusion Mamba (Fus-Mamba) block. Specifically, to solve the feature redundancy in multiple canning routes, the MSpa-Mamba block incorporates the multi-scale strategy to minimize the computational redundancy and alleviate the feature redundancy of SSM. In addition, Spe-Mamba is designed for spectral feature exploration, which is essential for HSI feature modeling. Moreover, to alleviate the heterogeneous gap between HSI and LiDAR/SAR data, we design Fus-Mamba block for multi-source feature fusion. The original Mamba is extended to accommodate dual inputs, and cross-modal feature interaction is enhanced. Extensive experimental results on three multi-source remote sensing datasets demonstrate the superiority performance of the proposed MSFMamba over the state-of-the-art models. Source codes of MSFMamba will be made public available at https://github.com/summitgao/MSFMamba.
@@
CNNs, Transformer models, Selective Structured State Space Model, Hyperspectral imaging, LiDAR/SAR data
@@ I Introduction
WITH the rapid development of remote sensing technology and sensor platforms, multi-source remote sensing data from various platforms such as satellites, aircraft, and drones have become increasingly abundant [1]. These data have been widely applied in land cover classification [2], geological resource exploration [3], urban planning [4], disaster warning [5], and agricultural monitoring [6]. Among these applications, land cover classification is an important and fundamental task [7] [8].
@@
Among these multi-source data captured by various sensors, hyperspectral image (HSI), light detection and ranging (LiDAR) data, and synthetic aperture radar (SAR) data are of great significance. HSI can identify subtle differences among ground objects through its rich spectral information [9] [10] [11]. HSI classification has been widely studied, since the rich spectral information contained in HSI presents great opportunities and challenges. On the other hand, LiDAR data precisely measure the three-dimensional structure of terrain and objects with high spatial resolution and accuracy [12] [13]. SAR sensors perform imaging in all weather conditions and at all times, penetrating clouds and vegetation [14] [15] [16]. The fusion of HSI and LiDAR/SAR data can leverage the complementary advantages of both data sources and lead to improved ground object classification performance [17] [18] [19]. By combining the spectral information of HSI with the three-dimensional structural information of LiDAR or the all weather imaging capability of SAR, the limitations of a single data source can be alleviated and more reliable classification results can be achieved [20] [21]. Therefore, in this paper, we mainly focus on HSI and LiDAR/SAR data joint classification.
@@
Because of diverse distributions and feature representations in multi-source images, the critical challenge for HSI and LiDAR/SAR data classification lies in how to bridge the heterogeneity gap. Recently, tremendous efforts have been made for multi-source data classification [22] [23]. The most common framework that demonstrated promising classification results employs CNN-based feature encoder in an end-to-end manner. Cross-guided attention [24], cross-channel correlation [25], federated learning [26] and cross-scale mixing attention [27] are employed for multi-source data classification. In addition, Transformer-based models have demonstrated outstanding long-range feature modeling capabilities in multi-source remote sensing data classification [28]. Spatial-spectral attention [29], hierarchical attention [30] [31], scale-adaptive attention [32] are used for HSI and LiDAR/SAR data classification.
Recently, an improved state-space model (SSM) [33] with a selective scanning mechanism, Mamba [34], has emerged as an effective alternative for computer vision and remote sensing image interpretation tasks. Compared with Transformer-based models, state-space sequence models reformulate the attention mechanism so that it scales linearly with the sequence length, and it significantly reduces the computational costs when the sequence is very long [35]. However, the SSM has rarely been explored to solve the multi-source remote sensing image classification task.
@@
It is a non-trivial and challenging task to introduce the SSM into multi-source data classification, due to the following reasons: 1) Feature redundancy in multiple scanning routes. To address the challenge of image data nondirectionality, existing SSM-based models generally use multi-scan strategy (Forward, backward, horizontal, and vertical scan) to ensure that every part of the image can establish connections with other parts. However, the multi-scan strategy significantly increases the feature redundancy. Therefore, how to reduce the information redundancy of SSM is one of the primary challenges we face. 2) The Heterogeneous gap between multi-source data. SSM’s role in multi-source remote sensing data classification has not yet been fully explored, as SSM lacks a design similar to cross-attention. This prompts us to investigate how to use SSM to alleviate the heterogeneous gap between HSI and LiDAR/SAR data.
@@
To address the above challenges, we propose Multi-Scale Feature Fusion Mamba (MSFMamba) network for the joint classification of HSI and LiDAR/SAR data. MSFMamba mainly comprises three parts: Multi-Scale Spatial Mamba (MSpa-Mamba) block, Spectral Mamba (Spe-Mamba) block, and Fusion Mamba (Fus-Mamba) block. Specifically, to solve the feature redundancy in multiple scanning routes, the MSpa-Mamba block incorporates the multi-scale strategy to minimize the computational redundancy and alleviate the feature redundancy in SSM. In addition, Spe-Mamba is designed for spectral feature exploration, which is essential for HSI feature modeling. Moreover, to alleviate the heterogeneous gap between HSI and LiDAR/SAR data, we design Fus-Mamba block for multi-source feature fusion. The original Mamba is extended to accommodate dual inputs, and cross-modal feature interaction is enhanced. Experimental results on three multi-source remote sensing datasets demonstrate the superior performance of the proposed MSFMamba.
@@
  • Multi-Scale Spatial Mamba (MSpa-Mamba) block is designed to incorporate the multi-scale strategy to minimize the computational burden. To the best of our knowledge, this is the first work to investigate the multi-scale SSMs for multi-source remote sensing image classification.
  • Fusion Mamba (Fus-Mamba) block is proposed to enhance cross-modal feature interactions between multi-source data.
  • Extensive experiments on three public benchmark datasets reveal that our proposed method outperforms the state-of-the-art methods. As a side contribution, we will release our codes to benefit the other researchers interested in remote sensing image interpretation.
The remainder of this paper is organized as follows: Section II reviews related works. Section III introduces the preliminaries of the state space model and the detailed structure of MSFMamba. Experiments on three multi-source datasets are presented in Section IV. Conclusion and future works are drawn in Section V.
@@ II-A Deep Learning-Based Multi-Source Image Classification
Deep learning-based multi-source data classification has witnessed significant progress recently. Convolutional neural networks (CNNs) have been widely used in an end-to-end manner for multi-source feature fusion. Dong et al. [24] used self-attention and cross-guided attention to emphasize interesting regions and channels of multi-source features. Gao et al. [27] presented a cross-mixing attention network to extract multi-scale features for hyperspectral and multispectral image joint classification. Xu et al. [36] introduced a dual-tunnel CNN architecture that separately extracts spectral-spatial features from HSI and LiDAR data. Zhang et al. [37] proposed a model known as patch-to-patch CNN for multi-source data classification. They constructed a three-tower structure to compute multi-scale features. Mohla et al. [38] proposed FusAtNet, which leverages HSI to generate an attention map via self-attention mechanism, which effectively highlights the intrinsic spectral features. In addition, cross-attention is employed to fuse the multi-source features. Fang et al. [39] presented the spatial–spectral enhancement module to promote multi-modal feature interaction.
To enhance the long-range feature dependencies, many Transformer-based methods have been proposed for multi-source remote sensing image classification. Hu et al. [40] proposed a simple yet effective parallel Transformer model. One Transformer acts as HSI feature encoder, and the other one is responsible for cross-modal feature interactions. Zhao et al. [41] proposed a CNN and Transformer hybrid model for HSI and LiDAR data classification. Cross-token attention is used for cross-modal feature fusion. Yao et al. [42] proposed Extended Vision Transformer (ExViT), which modifies the traditional ViT to handle land use and land cover classification tasks. This framework processes multimodal remote sensing image data through parallel branches and introduces a cross-modality attention module to enhance information exchange between different modalities. Li et al. [43] introduced mixing attention and convolution network for multi-source feature fusion. Although these CNN and Transformer-based methods have gained performance improvement, there is still an urgent need for an efficient and effective scheme to model the heterogeneous representation. In this paper, we explore Mamba’s potential for multi-source remote sensing feature fusion, providing an effective paradigm to propel the progress of heterogeneous feature fusion.
@@ II-B State Space Model
State-Space Models (SSM) [33] [44] [45] have become practical components for constructing deep networks due to their cutting-edge performance in analyzing continuous long sequential data. Gu et al. [45] introduced a diagonal structure and combined it with a diagonal plus low-rank approach to construct structured SSM. Smith et al. [33] improved the efficiency of SSM by introducing parallel scanning techniques. Mamba [34] incorporates data-dependent parameters to amend the linear time invariant characteristics of SSM-based models, and it has demonstrated excellent performance over Transformers on large-scale datasets.
Recently, inspired by pioneering SSMs, Mamba-like frameworks have been employed in computer vision and remote sensing image interpretation tasks, such as VMamba [46], RS3Mamba [47], and RSMamba [48]. RSMamba [48] employed dynamic multi-path activation mechanism to enhance the performance of remote sensing image classification. Zhang et al. [49] employed a unique encoder architecture, based on the Mamba design, to effectively extract semantic information from remote sensing images. Chen et al. [50] used a Mamba-like feature encoder to learn global spatial contextual information for remote sensing image change detection. Li et al. [51] proposed a spectral Mamba block to extract the spectral features of HSI, and improve the HSI classification performance. These works show the potential for extension into multi-source data fusion, its particular application in HSI and LiDAR/SAR data joint classification tasks remains unexplored.
@@ III-A Preliminaries
State-Space Model (SSM). SSM is a linear time-invariant system that map an input sequence $x(t)\in\mathbb{R}^{N}$ to an output sequence $y(t)\in\mathbb{R}^{N}$. They are mathematically represented by linear ordinary differential equations:
@@
where $h(t)\in\mathbb{R}^{N}$ indicate hidden state. $h^{\prime}(t)\in\mathbb{R}^{N}$ refers to the time derivative of $h(t)$. $N$ represents the number of states. Additionally, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the state transition matrix. $\mathbf{B}\in\mathbb{R}^{N\times 1}$ and $\mathbf{C}\in\mathbb{R}^{N\times 1}$ are projection matrices. $\mathbf{D}\in\mathbb{R}^{N\times 1}$ is served as a residual connected operation.
SSMs are continuous-time models, and pose challenges when incorporated into deep learning networks. To solve the problem, discrete versions of SSMs are proposed, and the ordinary differential equations can be discretized by the zeroth-order hold rule. A timescale parameter $\Delta$ is incorporated to convert the continuous parameters $\bm{\mathrm{A}}$, $\bm{\mathrm{B}}$ into discrete parameters $\bm{\overline{\mathrm{A}}}$, $\bm{\overline{\mathrm{B}}}$ as follows:
@@
$\bm{\overline{\mathrm{B}}}=(\Delta\bm{\mathrm{A}})^{-1}(\mathrm{exp}({\Delta\bm{\mathrm{A}}})-\mathrm{\bm{I}})\cdot\Delta\bm{\mathrm{B}}.$ (4)
The projection matrix $\overline{\mathbf{B}}$ can be approximated using the first-order Taylor series. After discretization, the ordinary differential equations of SSMs can be represented as follows:
@@ III-B Overall Framework of the Proposed MSFMamba
As depicted in Fig. 1, our MSFMamba consists of two parts: 1) The main network with a series of $L$ spatial-spectral Mamba modules to learn high-quality multi-source feature representations. 2) A classifier composed of two fully-connected layers for land cover classification. The number of spatial-spectral Mamba module $L$ is a critical parameter that will be discussed in the experiments.
@@
The Spatial-Spectral Mamba Module is the critical component in our MSFMamba. Details of the Spatial-Spectral Mamba Module are illustrated in Fig. 1(a). The input HSI and LiDAR/SAR features are first fed into the Multi-Scale Spatial Mamba (MSSpa-Mamba) block for spatial feature extraction. To further exploit the spectral features of HSI, the Spectral Mamba (Spe-Mamba) block is employed for HSI feature extraction. Then, we obtain optimized features $\mathbf{F}_{h}$ and $\mathbf{F}_{x}$. Next, $\mathbf{F}_{h}$ and $\mathbf{F}_{x}$ are fed into the Fusion Mamba block for cross-modal feature fusion and refinement. Finally, the refined HSI feature $\mathbf{F}^{\prime}_{h}$ and LiDAR/SAR feature $\mathbf{F}^{\prime}_{x}$ are generated.
As can be observed from Fig. 1(a), the MSSpa-Mamba block, Spe-Mamba block, and Fusion Mamba block are three key components in the Spatial-Spectral Mamba Module. We will give detailed descriptions of the three blocks in the following subsections.
III-C Multi-Scale Spatial Mamba Block (MSpa-Mamba)
Overview of the MSpa-Mamba. The Mamba can model long-range feature dependencies, which is critical for understanding the global context in remote sensing images. Existing SSM-based models commonly use multi-scan strategy to ensure that every part of the image can establish connections with other parts. However, the multi-scale strategy significantly increases the feature redundancy of SSM.
To solve the problem, we design a simple yet effective Multi-scale Spatial Mamba (MSpa-Mamba) block. As shown in Fig. 1(b), the input feature $\mathbf{X}$in $\mathbb{R}^{H\times W\times C}$ is fed into two parallel branches. In the first branch, the feature is processed by a linear layer, followed by a depth-wise convolution (DWConv), SiLU activation function, together with the Multi-scale Spatial SSM (MSpa-SSM) layer and LayerNorm. In the second branch, the feature is processed by a linear layer followed by the SiLU activation function. After that, features from the two branches are aggregated with element-wise multiplication. The computation of MSpa-Mamba block is as follows:

### v2

@@ Abstract
In the field of multi-source remote sensing image classification, remarkable progress has been made by using Convolutional Neural Network (CNN) and Transformer. While CNNs are constrained by their local receptive fields, Transformers mitigate this issue with their global attention mechanism. However, Transformers come with the trade-off of higher computational complexity. Recently, Mamba-based methods built upon the State Space Model (SSM) have shown great potential for long-range dependency modeling with linear complexity, but they have rarely been explored for multi-source remote sensing image classification tasks. To address this issue, we propose the Multi-Scale Feature Fusion Mamba (MSFMamba) network, a novel framework designed for the joint classification of hyperspectral image (HSI) and Light Detection and Ranging (LiDAR)/Synthetic Aperture Radar (SAR) data. The MSFMamba network is composed of three key components: the Multi-Scale Spatial Mamba (MSpa-Mamba) block, the Spectral Mamba (Spe-Mamba) block, and the Fusion Mamba (Fus-Mamba) block. The MSpa-Mamba block employs a multi-scale strategy to reduce computational cost and alleviate feature redundancy in multiple scanning routes, ensuring efficient spatial feature modeling. The Spe-Mamba block focuses on spectral feature extraction, addressing the unique challenges of HSI data representation. Finally, the Fus-Mamba block bridges the heterogeneous gap between HSI and LiDAR/SAR data by extending the original Mamba architecture to accommodate dual inputs, enhancing cross-modal feature interactions and enabling seamless data fusion. Together, these components enable MSFMamba to effectively tackle the challenges of multi-source data classification, delivering improved performance with optimized computational efficiency. Comprehensive experiments on four real-world multi-source remote sensing datasets (Berlin, Augsburg, Houston2018, and Houston2013) demonstrate the superiority of MSFMamba outperforms several state-of-the-art methods and achieves overall accuracies of 76.92%, 91.38%, 92.38%, and 92.86%, respectively. The source codes of MSFMamba will be publicly available at https://github.com/oucailab/MSFMamba.
@@
Transformer, state space model, hyperspectral image, synthetic aperture radar, light detection and ranging, multi-source data classification.
Figure 1: The effective receptive field of three models: (a) the CNN-based FusAtNet, (b) the Transformer-based ExViT, and (c) the proposed MSFMamba. A broader distribution of bright areas indicates larger effective receptive filed. The proposed MSFMamba achieves the largest effective receptive field.
@@ I Introduction
WITH the rapid development of remote sensing technology and sensor platforms, multi-source remote sensing data from various platforms such as satellites, aircraft, and drones have become increasingly abundant [1, 2, 3].These data have been widely applied in land use and land cover classification [4], geological resource exploration [5], urban planning [6], disaster warning [7], and agricultural monitoring [8] [9]. Among these applications, land use and land cover classification is an important and fundamental task [10] [11].
@@
Among these multi-source data captured by various sensors, HSI, LiDAR data, and SAR data are of great significance[12]. HSI can identify subtle differences among ground objects through its rich spectral information [13, 14, 15]. HSI classification has been widely studied, since the rich spectral information contained in HSI presents great opportunities and challenges. On the other hand, LiDAR data precisely measure the three-dimensional structure of terrain and objects with high spatial resolution and accuracy [16] [17]. SAR sensors perform imaging in all weather conditions and at all times, penetrating clouds and vegetation [18, 19, 20]. The fusion of HSI and LiDAR/SAR data can leverage the complementary advantages of multiple data sources and lead to improved ground object classification performance [21, 22, 23, 24]. By combining the spectral information of HSI with the three-dimensional structural information of LiDAR or the all weather imaging capability of SAR, the limitations of a single data source can be alleviated and more reliable classification results can be achieved [25] [26]. Therefore, in this paper, we mainly focus on HSI and LiDAR/SAR data joint classification.
@@
Because of diverse distributions and feature representations in multi-source images, the critical challenge for HSI and LiDAR/SAR data classification lies in how to bridge the heterogeneity gap. Recently, tremendous efforts have been made for multi-source data classification [27, 28, 29, 30].The most common framework that demonstrated promising classification results employs CNN-based feature encoder in an end-to-end manner. Cross-guided attention [31], cross-channel correlation [32], federated learning [33] and cross-scale mixing attention [34] are employed for multi-source data classification. CNN-based methods design complex attention mechanisms to capture useful features, but due to the inherent limitations of convolutional units, their receptive field range is limited, making it difficult to capture long-range dependencies. In addition, Transformer-based models have demonstrated outstanding long-range feature modeling capabilities in multi-source remote sensing data classification [35]. Spatial-spectral attention [36], hierarchical attention [37] [38], scale-adaptive attention [39] are used for HSI and LiDAR/SAR data classification. Although the Transformer has a global receptive field, its attention mechanism has high computational complexity and is less efficient when handling long-range dependencies.
Recently, an improved SSM [40] with a selective scanning mechanism, Mamba [41], has emerged as an effective alternative for computer vision and remote sensing image interpretation. As shown in Fig. 1, the Mamba-based method[42][43] performs better than CNN and Transformer-based methods in effective receptive field modeling. By expanding the receptive field, the model is able to capture broader spatial dependencies, which is crucial for recognizing contextual objects and larger patterns in remote sensing data. In contrast to Transformer-based models, the state-space sequence model reformulates the attention mechanism to scale linearly with sequence length, significantly reducing computational costs, especially for long sequences [44].
@@
It is a non-trivial and challenging task to introduce the SSM into multi-source data classification, due to the following reasons: 1) Feature redundancy in multiple scanning routes. To address the challenge of image data nondirectionality, existing SSM-based models generally use a multi-scan strategy (e.g., forward, backward, horizontal, and vertical scan) to ensure that every part of the image can establish connections with other parts. However, the multi-scan strategy significantly increases feature redundancy, as similar patterns are repeatedly extracted across overlapping scanning routes. Multi-scale feature extraction provides a promising solution to this challenge by capturing spatial dependencies at different levels of granularity. By combining fine-grained local details with broader contextual information, the model can reduce redundancy while preserving the most informative features. Thus, exploring an effective multi-scale strategy is critical for improving the efficiency of SSM-based approaches. 2) The Heterogeneous gap between multi-source data. SSM’s role in multi-source remote sensing data classification has not yet been fully explored, as SSM lacks a design similar to cross-attention. This prompts us to investigate how to use SSM to alleviate the heterogeneous gap between HSI and LiDAR/SAR data.
@@
To address the above challenges, we propose Multi-Scale Feature Fusion Mamba (MSFMamba) network for the joint classification of HSI and LiDAR/SAR data. MSFMamba mainly comprises three parts: MSpa-Mamba block, Spe-Mamba block, and Fus-Mamba block. Specifically, to solve the feature redundancy in multiple scanning routes, the MSpa-Mamba block incorporates the multi-scale strategy to minimize the computational redundancy and alleviate the feature redundancy in SSM. In addition, Spe-Mamba is designed for spectral feature exploration, which is essential for HSI feature modeling. Moreover, to alleviate the heterogeneous gap between HSI and LiDAR/SAR data, we design Fus-Mamba block for multi-source feature fusion. The original Mamba is extended to accommodate dual inputs, and cross-modal feature interaction is enhanced.
@@
  • We propose MSFMamba, a simple yet effective sequential scanning model for HSI and LiDAR/SAR data interpretation, which investigates the use of SSMs for multi-source remote sensing image classification.
  • We design a MSpa-Mamba block that incorporates a multi-scale strategy to minimize computational burden, along with a Fus-Mamba block to enhance cross-modal feature interactions between multi-source data.
  • Extensive experiments on four public benchmark datasets demonstrate that our method outperforms state-of-the-art techniques in multi-source remote sensing image classification.
The remainder of this paper is organized as follows. Section II reviews related work. Section III introduces the preliminaries of the state space model and the detailed structure of MSFMamba. Experiments on four multi-source datasets are presented in Section IV. Conclusions are drawn, and future work is discussed in Section V.
@@ II-A Deep Learning-Based Multi-Source Image Classification
Deep learning-based remote sensing image analysis has witnessed significant progress recently [45, 46, 47, 48]. Convolutional neural networks (CNNs) have been widely used in an end-to-end manner for multi-source feature fusion. Dong et al. [31] used self-attention and cross-guided attention to emphasize regions and channels of interest. Gao et al. [34] presented a cross-mixing attention network to extract multi-scale features for hyperspectral and multispectral image joint classification. Xu et al. [49] introduced a dual-tunnel CNN architecture that separately extracts spectral-spatial features from HSI and LiDAR data. Zhang et al. [50] proposed a model known as patch-to-patch CNN for multi-source data classification. They constructed a three-tower structure to compute multi-scale features. Mohla et al. [51] proposed Feature Fusion And exTraction Network (FusAtNet), which leverages HSI to generate an attention map via self-attention mechanism, which effectively highlights the intrinsic spectral features. In addition, cross-attention is employed to fuse the multi-source features. Fang et al. [52] presented a spatial–spectral enhancement module to promote multi-modal feature interaction. In addition,
To enhance the long-range feature dependencies, many Transformer-based methods have been proposed for multi-source remote sensing image classification. Hu et al. [53] proposed a simple yet effective parallel Transformer model. One Transformer acts as HSI feature encoder, and the other one is responsible for cross-modal feature interactions. Zhao et al. [54] proposed a CNN and Transformer hybrid model for HSI and LiDAR data classification. Cross-token attention is used for cross-modal feature fusion. Yao et al. [55] proposed Extended Vision Transformer (ExViT), which modifies the traditional Vision Transformer to handle land use and land cover classification tasks. This framework processes multimodal remote sensing image data through parallel branches and introduces a cross-modality attention module to enhance information exchange between different modalities. Li et al. [56] introduced mixing attention and convolution network for multi-source feature fusion. Although these CNN and Transformer-based methods have gained performance improvement, there is still an urgent need for an efficient and effective scheme for heterogeneous representation. In this paper, we explore Mamba’s potential for multi-source remote sensing feature fusion, and provide an effective paradigm to propel the progress of heterogeneous feature fusion.
@@ II-B State Space Model
SSMs [57][58] have become practical components for constructing deep networks due to their cutting-edge performance in analyzing continuous long sequential data. Gu et al. [58] introduced a diagonal structure and combined it with a diagonal plus low-rank approach to construct structured SSM. Smith et al. [40] improved the efficiency of SSM by introducing parallel scanning techniques. Mamba [41] incorporates data-dependent parameters to amend the linear time invariant characteristics of SSM-based models, and it has demonstrated excellent performance over Transformers on large-scale datasets.
Recently, inspired by pioneering SSMs, Mamba-like frameworks have been employed in computer vision and remote sensing image interpretation tasks, such as Visual State Space Model (VMamba) [59], remote sensing images semantic segmentation Mamba (RS3Mamba) [60], and Remote Sensing Mamba(RSMamba) [61]. RSMamba [61] employed dynamic multi-path activation mechanism to enhance the performance of remote sensing image classification. Zhang et al. [62] employed a unique encoder architecture, based on the Mamba design, to effectively extract semantic information from remote sensing images. Chen et al. [63] used a Mamba-like feature encoder to learn global spatial contextual information for remote sensing image change detection. Li et al. [64] proposed a spectral Mamba block to extract the spectral features of HSI, and improve the HSI classification performance. These works show the potential for Mamba’s extension into multi-source data fusion, but its particular application in the joint classification task of HSI and LiDAR/SAR data remains unexplored.
@@ III-A Preliminaries
State-Space Model. SSM is a linear time-invariant system that maps an input sequence $x(t)\in\mathbb{R}^{N}$ to an output sequence $y(t)\in\mathbb{R}^{N}$. They are mathematically represented by linear ordinary differential equations:
@@
where $h(t)\in\mathbb{R}^{N}$ indicates a hidden state, $h^{\prime}(t)\in\mathbb{R}^{N}$ refers to the time derivative of $h(t)$, and $N$ represents the number of states. Additionally, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the state transition matrix, $\mathbf{B}\in\mathbb{R}^{N\times 1}$ , $\mathbf{C}\in\mathbb{R}^{N\times 1}$ are projection matrices, and $\mathbf{D}\in\mathbb{R}^{N\times 1}$ is served as a residual connected operation.
SSMs are continuous-time models, and it is challenging for them to be incorporated into deep learning networks. To solve the problem, discrete versions of SSMs are proposed, and the ordinary differential equations can be discretized by the zeroth-order hold rule. A timescale parameter $\Delta$ is incorporated to convert the continuous parameters $\bm{\mathrm{A}}$, $\bm{\mathrm{B}}$ into discrete parameters $\bm{\overline{\mathrm{A}}}$, $\bm{\overline{\mathrm{B}}}$ , respectively, as:
@@
$\bm{\overline{\mathrm{B}}}=(\Delta\bm{\mathrm{A}})^{-1}(\mathrm{exp}({\Delta\bm{\mathrm{A}}})-\mathrm{\bm{I}})\cdot\Delta\bm{\mathrm{B}}\approx\Delta\bm{\mathrm{B}}.$ (4)
In practice, as noted in [41], the projection matrix $\overline{\mathbf{B}}$ can be approximated by applying a first-order Taylor expansion to the term involving the matrix exponential. After discretization, the ordinary differential equations of SSMs can be represented as follows:
@@ III-B Overall Framework of the Proposed MSFMamba
As depicted in Fig. 2, our MSFMamba consists of two parts: 1) The main network with a series of $L$ spatial-spectral Mamba modules to learn high-quality multi-source feature representations. 2) A classifier with two fully-connected layers for land cover classification. The number of spatial-spectral Mamba module $L$ is a critical parameter that will be discussed in the experiments.
@@
The Spatial-Spectral Mamba Module is the critical component in our MSFMamba. Details of the Spatial-Spectral Mamba Module are shown in Fig. 2(a). The input HSI and LiDAR/SAR features are first fed into the MSpa-Mamba block for spatial feature extraction. To further exploit the spectral features of HSI, the Spe-Mamba block is employed for HSI feature extraction. Then, we obtain optimized features $\mathbf{F}_{h}$ and $\mathbf{F}_{x}$. Next, $\mathbf{F}_{h}$ and $\mathbf{F}_{x}$ are fed into the Fusion Mamba block for cross-modal feature fusion and refinement. Finally, the refined HSI feature $\mathbf{F}^{\prime}_{h}$ and LiDAR/SAR feature $\mathbf{F}^{\prime}_{x}$ are generated.
As can be observed from Fig. 2(a), the MSpa-Mamba block, Spe-Mamba block, and Fusion Mamba block are three key components in the Spatial-Spectral Mamba Module. We will give detailed descriptions of the three blocks in the following subsections.
III-C MSpa-Mamba Block
Overview of the MSpa-Mamba. The Mamba can model long-range feature dependencies, which is critical for understanding the global context in remote sensing images. Existing SSM-based models commonly use multi-scan strategy to ensure that every part of the image can establish connections with other parts. However, the multi-scan strategy significantly increases the feature redundancy of SSM.
To solve the problem, we design a simple yet effective MSpa-Mamba block. As shown in Fig. 2(b), the input feature $\mathbf{X}$in $\mathbb{R}^{H\times W\times C}$ is fed into two parallel branches. In the first branch, the feature is processed by a linear layer, followed by a depth-wise convolution (DWConv), Sigmoid Linear Unit (SiLU) activation function, together with the Multi-scale Spatial SSM (MSpa-SSM) layer and LayerNorm. In the second branch, the feature is processed by a linear layer followed by the SiLU activation function. After that, features from the two branches are aggregated with element-wise multiplication. The computation of MSpa-Mamba block is as follows: